# Antenna Position Optimization for Movable Antenna-Empowered Near-Field Sensing

### v3

@@ I Introduction
Notations: Boldface lower and upper cases are used to represent vectors and matrices, respectively. The conjugate of a complex is denoted by $(\cdot)^{*}$. The transpose and conjugate transpose of a matrix/vector are denoted by $(\cdot)^{\top}$ and $(\cdot)^{\mathrm{H}}$, respectively. The sets of $N_{1}\times N_{2}$ dimensional real and complex matrices are denoted by $\mathbb{R}^{N_{1}\times N_{2}}$ and $\mathbb{C}^{N_{1}\times N_{2}}$, respectively. The $l_{2}$-norm of vector $\boldsymbol{v}$ is denoted by $\lVert\boldsymbol{v}\rVert$. For a real number $c$, $\lvert c\rvert$ denotes its absolute value and $\lfloor c\rfloor$ denotes the maximum integer that is no larger than $c$. The statistical expectation is denoted by $\mathbb{E}\{\cdot\}$. $\boldsymbol{I}_{N}$ denotes the $N$-dimensional identity matrix.
@@ III-A AoA Estimation
$\text{CRB}_{u}(u;\boldsymbol{x})=\frac{\sigma^{2}\lambda^{2}}{8\pi^{2}TPN\lvert\beta\rvert^{2}}\cdot\frac{1}{F_{u}(\boldsymbol{x},u)}\leq\text{MSE}(u),$ (10)
where
@@ III-B Problem Formulation and Proposed Solution
Our objective is to minimize $\text{CRB}_{u}(u;\boldsymbol{x})$ by optimizing the APV $\boldsymbol{x}$. However, the CRB in (10) is dependent on both the APV $\boldsymbol{x}$ and AoA itself. To tackle this issue, we focus on minimizing the worst-case $\text{CRB}_{u}(u;\boldsymbol{x})$ for all possible values of $u$, i.e., $\max_{u}\text{CRB}_{u}(u;\boldsymbol{x})$. The associated min-max problem can be easily shown equivalent to the following max-min problem based on (10), i.e.,
$\min_{\boldsymbol{x}}\max_{u\in[0,1)}\text{CRB}_{u}(u;\boldsymbol{x})\iff\max_{\boldsymbol{x}}\min_{u\in[0,1)}F_{u}(\boldsymbol{x},u).$ (12)
@@
where $d$ denotes the minimum inter-MA distance to avoid mutual coupling, and $u_{\text{opt}}$ is the AoA value that yields the worst-case CRB on the AoA, i.e., $u_{\text{opt}}=\arg\max_{u}\text{CRB}_{u}(u;\boldsymbol{x})$. Note that since $u_{\text{opt}}\geq 0$, it is desirable that the MAs should be positioned as dispersive apart as possible to maximize $F_{u}^{\star}(\boldsymbol{x})$, which helps increase the variance terms $\text{var}(\boldsymbol{x})$ and $\text{var}(\tilde{\boldsymbol{x}})$, as well as the covariance term $\text{cov}(\boldsymbol{x},\tilde{\boldsymbol{x}})$. This will be rigorously proved in Theorem 1 below.
@@ IV Antenna Position Optimization for Distance Estimation
In this section, we consider Case 2 where the spatial AoA is already known, denoted as $u^{\star}$. To estimate the distance $r$, we also apply the MUSIC algorithm by leveraging the distance-related information in the signal phase [18]. For simplicity, the detailed process of adopting the MUSIC algorithm for the distance estimation is omitted here. The associated MSE and CRB of the distance estimation are given by
$\text{MSE}(r)\geq\text{CRB}_{r}(r;\boldsymbol{x})=\sigma^{2}\lambda^{2}\big{/}\big{(}8\pi^{2}TPN\lvert\beta\rvert^{2}\big{)}\cdot F_{r}^{-1}(\boldsymbol{x},r),$ (15)
@@
Note that the CRB in (15) depends on the distance $r$. To eliminate its effects, similar to the AoA estimation, we aim to minimize the maximum of the right-hand side of (15) among all possible values of $r$, i.e., $\max_{r}\text{CRB}_{r}(r;\boldsymbol{x})$. Hence, the corresponding min-max problem can be reformulated as a max-min problem, i.e.,
$\min_{\boldsymbol{x}}\max_{r\in[r_{\text{min}},r_{\text{max}}]}\text{CRB}_{r}(r;\boldsymbol{x})\iff\max_{\boldsymbol{x}}\min_{r\in[r_{\text{min}},r_{\text{max}}]}F_{r}(\boldsymbol{x},r).$ (17)
@@
where $r_{\text{opt}}$ is the distance value that yields the worst-case CRB on the distance, i.e., $r_{\text{opt}}=\arg\max_{r}\text{CRB}_{r}(r;\boldsymbol{x})$. It can be easily proved that the optimal APV to (P2) the same as that provided in Theorem 1 via a similar process to Appendix A in [13]. As a result, in the individual estimation of $u$ or $r$, the associated optimal APVs are identical and the same as the AoA estimation in the far-field sensing. The optimal MA positions should maximally increase the aperture to ensure the sensing resolution in the angle or distance domain.
@@ V Antenna Position Optimization for Joint AoA and Distance Estimation
$\text{FIM}(\boldsymbol{\eta})=\begin{bmatrix}J_{uu}&J_{ur}\\ J_{ru}&J_{rr}\end{bmatrix},$ (20)
where the detailed derivations of the FIMs $J_{\alpha\beta}(\alpha,\beta\in\{u,r\})$ are given in Appendix. Then, the CRB matrix is derived by taking the inverse of the FIM of the estimator, i.e., $\text{CRB}_{\boldsymbol{\eta}}(\boldsymbol{\eta};\boldsymbol{x})=\text{FIM}^{-1}(\boldsymbol{\eta})$. Specifically, the CRBs on AoA and distance in the joint estimation are given by
$\text{CRB}_{u}(\boldsymbol{x})=\text{var}(\boldsymbol{\tilde{x}})\big{/}\big{(}\text{var}(\boldsymbol{x})\text{var}(\boldsymbol{\tilde{x}})-{\text{cov}^{2}(\boldsymbol{x},\boldsymbol{\tilde{x}})}\big{)},$ (21)
$\text{CRB}_{r}(u,r;\boldsymbol{x})=\frac{4r^{4}\text{var}(\boldsymbol{x})+8ur^{3}{\text{cov}(\boldsymbol{x},\boldsymbol{\tilde{x}})}+4u^{2}r^{2}\text{var}(\boldsymbol{\tilde{x}})}{(1-u^{2})^{2}\Big{(}\text{var}(\boldsymbol{x})\text{var}(\boldsymbol{\tilde{x}})-\text{cov}^{2}(\boldsymbol{x},\boldsymbol{\tilde{x}})\Big{)}},$ (22)
respectively. The detailed procedure for deriving the CRBs is given in Appendix as well. It can be observed from (21) and (22) that there may exist a fundamental trade-off between minimizing $\text{CRB}_{u}(\boldsymbol{x})$ and $\text{CRB}_{r}(u,r;\boldsymbol{x})$ due to the complicated coupling between the variance and covariance terms therein. Moreover, although $\text{CRB}_{u}(\boldsymbol{x})$ is independent of the spatial AoA $u$, $\text{CRB}_{r}(u,r;\boldsymbol{x})$ depends on both the spatial AoA $u$ and distance $r$. To overcome this difficulty, we aim to minimize the sum of $\text{CRB}_{u}(\boldsymbol{x})$ and the worst-case $\text{CRB}_{r}(u,r;\boldsymbol{x})$, i.e., $\text{CRB}_{u}(\boldsymbol{x})+\max_{u,r}\text{CRB}_{r}(u,r;\boldsymbol{x})$, by optimizing the APV $\boldsymbol{x}$. Denote $\boldsymbol{\eta}_{\text{opt}}$ as the estimator vector that yields the worst-case CRB on the distance. Then, the associated optimization problem can be equivalently transformed into the following problem, i.e.,
(P3) $\max_{\boldsymbol{x}}\quad F_{\boldsymbol{\eta}}^{\star}(\boldsymbol{x})\triangleq\text{tr}^{-1}\Big{(}\text{CRB}_{\boldsymbol{\eta}}(\boldsymbol{\eta};\boldsymbol{x})\Big{)}\Big{|}_{\boldsymbol{\eta}=\boldsymbol{\eta}_{\text{opt}}}$ (23a)
@@
Figure 3: The CRB of the AoA estimation MSE versus the received SNR.
Figure 4: The CRB of the distance estimation MSE versus the received SNR.
@@
Figure 6: The sum of the joint estimation CRBs versus the received SNR.
@@ VII Conclusion
$\text{FIM}(\boldsymbol{\eta})=\frac{2}{\sigma^{2}}\sum_{t=1}^{T}\Re\Bigg{\{}s_{t}^{*}\boldsymbol{\Psi}(\boldsymbol{x},\boldsymbol{\eta})^{\mathsf{H}}\bigg{(}\boldsymbol{I}_{N}-\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})\Big{(}\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})^{\mathsf{H}}\\ \boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})\Big{)}^{-1}\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})^{\mathsf{H}}\bigg{)}\boldsymbol{\Psi}(\boldsymbol{x},\boldsymbol{\eta})s_{t}\Bigg{\}},$ (26)
@@
Hence, by (26), the four elements of $\text{FIM}(\boldsymbol{\eta})$ are given by
$J_{uu}=\text{var}(\boldsymbol{x})+\frac{2u}{r}\text{cov}(\boldsymbol{x},\boldsymbol{\tilde{x}})+\frac{u^{2}}{r^{2}}\text{var}(\boldsymbol{\tilde{x}}),$ (28)
$J_{ur}=J_{ru}=\frac{1-u^{2}}{2r^{2}}\Big{(}\text{cov}(\boldsymbol{x},\boldsymbol{\tilde{x}})+\frac{u}{r}\text{var}(\boldsymbol{\tilde{x}})\Big{)},$ (29)
$J_{rr}=\Big{(}\frac{1-u^{2}}{2r^{2}}\Big{)}^{2}\text{var}(\boldsymbol{\tilde{x}}),$ (30)
respectively. Then, the CRB matrix for the joint estimation is obtained by taking the inverse of $\text{FIM}(\boldsymbol{\eta})$, i.e.,
$\text{CRB}_{\boldsymbol{\eta}}(\boldsymbol{\eta};\boldsymbol{x})=\begin{bmatrix}\Big{(}J_{uu}-\dfrac{J_{ur}J_{ru}}{J_{rr}}\Big{)}^{-1}&\Big{(}J_{ru}-\dfrac{J_{uu}J_{rr}}{J_{ur}}\Big{)}^{-1}\\[8.0pt] \Big{(}J_{ur}-\dfrac{J_{uu}J_{rr}}{J_{ru}}\Big{)}^{-1}&\Big{(}J_{rr}-\dfrac{J_{ru}J_{ur}}{J_{uu}}\Big{)}^{-1}\end{bmatrix}.$ (31)

### v4

@@ I Introduction
Notations: Boldface lower and upper cases are used to represent vectors and matrices, respectively. The conjugate of a complex is denoted by $(\cdot)^{*}$. The transpose and conjugate transpose of a matrix/vector are denoted by $(\cdot)^{\top}$ and $(\cdot)^{\mathrm{H}}$, respectively. The sets of $N_{1}\times N_{2}$ dimensional real and complex matrices are denoted by $\mathbb{R}^{N_{1}\times N_{2}}$ and $\mathbb{C}^{N_{1}\times N_{2}}$, respectively. The $l_{2}$-norm of vector $\boldsymbol{v}$ is denoted by $\lVert\boldsymbol{v}\rVert$. For a real number $c$, $\lvert c\rvert$ denotes its absolute value and $\lfloor c\rfloor$ denotes the maximum integer that is no larger than $c$. The expectation is denoted by $\mathbb{E}\{\cdot\}$. $\boldsymbol{I}_{N}$ denotes the $N$-dimensional identity matrix.
@@ III-A AoA Estimation
$\text{CRB}_{u}(\boldsymbol{x},u)=\frac{\kappa}{F_{u}(\boldsymbol{x},u)}\leq\text{MSE}(u),$ (10)
where $\kappa\triangleq\frac{\sigma^{2}\lambda^{2}}{8\pi^{2}TPN\lvert\beta\rvert^{2}}$ and
@@ III-B Problem Formulation and Proposed Solution
Our objective is to minimize $\text{CRB}_{u}(\boldsymbol{x},u)$ by optimizing the APV $\boldsymbol{x}$. However, the CRB in (10) is dependent on both the APV $\boldsymbol{x}$ and AoA itself. To tackle this issue, we focus on minimizing the worst-case $\text{CRB}_{u}(\boldsymbol{x},u)$ for all possible values of $u$, i.e., $\max_{u}\text{CRB}_{u}(\boldsymbol{x},u)$. The associated min-max problem can be easily shown equivalent to the following max-min problem based on (10), i.e.,
$\min_{\boldsymbol{x}}\max_{u\in[0,1)}\text{CRB}_{u}(\boldsymbol{x},u)\iff\max_{\boldsymbol{x}}\min_{u\in[0,1)}F_{u}(\boldsymbol{x},u).$ (12)
@@
where $d$ denotes the minimum inter-MA distance to avoid mutual coupling, and $u_{\text{opt}}$ is the AoA value that yields the worst-case CRB on the AoA, i.e., $u_{\text{opt}}=\arg\max_{u}\text{CRB}_{u}(\boldsymbol{x},u)$. Note that since $u_{\text{opt}}\geq 0$, it is desirable that the MAs should be positioned as dispersive apart as possible to maximize $F_{u}^{\star}(\boldsymbol{x})$, which helps increase the variance terms $\text{var}(\boldsymbol{x})$ and $\text{var}(\tilde{\boldsymbol{x}})$, as well as the covariance term $\text{cov}(\boldsymbol{x},\tilde{\boldsymbol{x}})$. This will be rigorously proved in Theorem 1 below.
@@ IV Antenna Position Optimization for Distance Estimation
In this section, we consider Case 2 where the spatial AoA is already known, denoted as $u^{\star}$. To estimate the distance $r$, we also apply the MUSIC algorithm by leveraging the distance-related information in the signal phase [18]. For simplicity, the detailed process for distance estimation is omitted here. The associated MSE and CRB are given by
$\text{MSE}(r)\geq\text{CRB}_{r}(\boldsymbol{x},r)=\kappa\cdot F_{r}^{-1}(\boldsymbol{x},r),$ (15)
@@
Note that the CRB in (15) depends on the distance $r$. To eliminate its effects, similar to the AoA estimation, we aim to minimize the maximum of the right-hand side of (15) among all possible values of $r$, i.e., $\max_{r}\text{CRB}_{r}(\boldsymbol{x},r)$. Hence, the corresponding min-max problem can be reformulated as a max-min problem, i.e.,
$\min_{\boldsymbol{x}}\max_{r\in[r_{\text{min}},r_{\text{max}}]}\text{CRB}_{r}(\boldsymbol{x},r)\iff\max_{\boldsymbol{x}}\min_{r\in[r_{\text{min}},r_{\text{max}}]}F_{r}(\boldsymbol{x},r).$ (17)
@@
where $r_{\text{opt}}$ is the distance value that yields the worst-case CRB on the distance, i.e., $r_{\text{opt}}=\arg\max_{r}\text{CRB}_{r}(\boldsymbol{x},r)$. It can be easily proved that the optimal APV to (P2) the same as that provided in Theorem 1 via a similar process to Appendix A in [13]. As a result, in the individual estimation of $u$ or $r$, the associated optimal APVs are identical and the same as the AoA estimation in the far-field sensing. The optimal MA positions should maximally increase the aperture to ensure the sensing resolution in the angle or distance domain.
@@ V Antenna Position Optimization for Joint AoA and Distance Estimation
$\text{FIM}(\boldsymbol{x},\boldsymbol{\eta})=\begin{bmatrix}J_{uu}&J_{ur}\\ J_{ru}&J_{rr}\end{bmatrix},$ (20)
where the detailed derivations of the FIMs $J_{\alpha\beta}(\alpha,\beta\in\{u,r\})$ are given in Appendix. Then, the CRB matrix is derived by taking the inverse of the FIM of the estimator, i.e., $\text{CRB}_{\boldsymbol{\eta}}(\boldsymbol{x},\boldsymbol{\eta})=\text{FIM}^{-1}(\boldsymbol{x},\boldsymbol{\eta})$. Specifically, the CRBs on AoA and distance in the joint estimation are given by
$\text{CRB}_{u}(\boldsymbol{x})=\kappa\cdot\text{var}(\boldsymbol{\tilde{x}})\big{/}\big{(}\text{var}(\boldsymbol{x})\text{var}(\boldsymbol{\tilde{x}})-{\text{cov}^{2}(\boldsymbol{x},\boldsymbol{\tilde{x}})}\big{)},$ (21)
$\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})=\frac{\kappa\Big{(}4r^{4}\text{var}(\boldsymbol{x})+8ur^{3}{\text{cov}(\boldsymbol{x},\boldsymbol{\tilde{x}})}+4u^{2}r^{2}\text{var}(\boldsymbol{\tilde{x}})\Big{)}}{(1-u^{2})^{2}\Big{(}\text{var}(\boldsymbol{x})\text{var}(\boldsymbol{\tilde{x}})-\text{cov}^{2}(\boldsymbol{x},\boldsymbol{\tilde{x}})\Big{)}},$ (22)
respectively. The detailed procedure for deriving the CRBs is given in Appendix as well. It can be observed from (21) and (22) that there may exist a fundamental trade-off between minimizing $\text{CRB}_{u}(\boldsymbol{x})$ and $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ due to the complicated coupling between the variance and covariance terms therein. Moreover, although $\text{CRB}_{u}(\boldsymbol{x})$ is independent of the spatial AoA $u$, $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$ depends on both the spatial AoA $u$ and distance $r$. To overcome this difficulty, we aim to minimize the sum of $\text{CRB}_{u}(\boldsymbol{x})$ and the worst-case $\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$, i.e., $\text{CRB}_{u}(\boldsymbol{x})+\max_{u,r}\text{CRB}_{r}(\boldsymbol{x},\boldsymbol{\eta})$, by optimizing the APV $\boldsymbol{x}$. Denote $\boldsymbol{\eta}_{\text{opt}}$ as the estimator vector that yields the worst-case CRB on the distance. Then, the associated optimization problem can be equivalently transformed into the following problem, i.e.,
(P3) $\max_{\boldsymbol{x}}\quad F_{\boldsymbol{\eta}}^{\star}(\boldsymbol{x})\triangleq\text{Tr}^{-1}\Big{(}\text{CRB}_{\boldsymbol{\eta}}(\boldsymbol{x},\boldsymbol{\eta})\Big{)}\Big{|}_{\boldsymbol{\eta}=\boldsymbol{\eta}_{\text{opt}}}$ (23a)
@@
Figure 3: CRB of the AoA estimation MSE versus the received SNR.
Figure 4: CRB of the distance estimation MSE versus the received SNR.
@@
Figure 6: Sum of the joint estimation CRBs versus the received SNR.
@@ VII Conclusion
$\text{FIM}(\boldsymbol{x},\boldsymbol{\eta})=\frac{2}{\sigma^{2}}\sum_{t=1}^{T}\Re\Bigg{\{}s_{t}^{*}\boldsymbol{\Psi}(\boldsymbol{x},\boldsymbol{\eta})^{\mathsf{H}}\bigg{(}\boldsymbol{I}_{N}-\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})\\ \Big{(}\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})\Big{)}^{-1}\boldsymbol{\alpha}(\boldsymbol{x},\boldsymbol{\eta})^{\mathsf{H}}\bigg{)}\boldsymbol{\Psi}(\boldsymbol{x},\boldsymbol{\eta})s_{t}\Bigg{\}},$ (26)
@@
Hence, by (26), the four elements of $\text{FIM}(\boldsymbol{x},\boldsymbol{\eta})$ are given by
$J_{uu}=\kappa^{-1}\cdot\Big{(}\text{var}(\boldsymbol{x})+\frac{2u}{r}\text{cov}(\boldsymbol{x},\boldsymbol{\tilde{x}})+\frac{u^{2}}{r^{2}}\text{var}(\boldsymbol{\tilde{x}})\Big{)},$ (28)
$J_{ur}=J_{ru}=\kappa^{-1}\cdot\frac{1-u^{2}}{2r^{2}}\Big{(}\text{cov}(\boldsymbol{x},\boldsymbol{\tilde{x}})+\frac{u}{r}\text{var}(\boldsymbol{\tilde{x}})\Big{)},$ (29)
$J_{rr}=\kappa^{-1}\cdot\Big{(}\frac{1-u^{2}}{2r^{2}}\Big{)}^{2}\text{var}(\boldsymbol{\tilde{x}}),$ (30)
respectively. Then, the CRB matrix for the joint estimation is obtained by taking the inverse of $\text{FIM}(\boldsymbol{x},\boldsymbol{\eta})$, i.e.,
$\text{CRB}_{\boldsymbol{\eta}}(\boldsymbol{x},\boldsymbol{\eta})=\begin{bmatrix}\Big{(}J_{uu}-\dfrac{J_{ur}J_{ru}}{J_{rr}}\Big{)}^{-1}&\Big{(}J_{ru}-\dfrac{J_{uu}J_{rr}}{J_{ur}}\Big{)}^{-1}\\[8.0pt] \Big{(}J_{ur}-\dfrac{J_{uu}J_{rr}}{J_{ru}}\Big{)}^{-1}&\Big{(}J_{rr}-\dfrac{J_{ru}J_{ur}}{J_{uu}}\Big{)}^{-1}\end{bmatrix}.$ (31)